OpenStudy (anonymous):

find the numbers of permutations 10P6

jimthompson5910 (jim_thompson5910):

do you have any ideas on the first step?

OpenStudy (anonymous):

\[_{10} P _{6}\]

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

Are you familiar with the permutation formula?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

If not, the permutation formula is n P r = (n!)/((n-r)!) where n! means n! = n*(n-1)*(n-2)*...*3*2*1 For example: 6! = 6*5*4*3*2*1

jimthompson5910 (jim_thompson5910):

So in our case, 10 P 6 is 10 P 6 = (10!)/((10-6)!) since n = 10 and r = 6

jimthompson5910 (jim_thompson5910):

Can you take it from here or do you need more help?

OpenStudy (anonymous):

yes i need more help please

jimthompson5910 (jim_thompson5910):

(10!)/((10-6)!) will become (10!)/(4!) since 10-6 = 4

jimthompson5910 (jim_thompson5910):

Then you expand out 10! like so 10*9*8*7*6*5*4! We stop at 4 because the "4!" terms will cancel So (10!)/(4!) turns into (10*9*8*7*6*5*4!)/(4!) and the "4!" terms will cancel leaving us with 10*9*8*7*6*5

jimthompson5910 (jim_thompson5910):

Last step is to multiply out 10*9*8*7*6*5 to get your answer